OpenStudy (anonymous):

There is a 2/3 chance of rain for each of three days. If the weather on each day is independent of the weather on the other two days, what is the probability that it will rain on none of the three days?

OpenStudy (kropot72):

The chance of it not raining for each of three days is 1 - (2/3) = 1/3 The probability that it will rain on none of the three days is therefore \[\frac{1}{3}\times\frac{1}{3}\times\frac{1}{3}=you\ can\ calculate\]